Shift report 20010817 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Aug 17 2001 - 16:46:30 EDT

  • Next message: Apache: "Shift report 20010818 16:00-24:00"

    ------------------------------------------------------------
    Date: Fri Aug 17 16:46:29 2001, Shift: 08:00-16:00
    Supervisor: JHL, with: 
    Log book#7 pages 109 - 116
    
    Shift Summary:
    ------------------------------------------------------------
    Machine/Beam
    - 3-4 Hours Cogged Beam in this shift
    - DAQ rates: 10-20/sec with triggers 1,4,8
    
    Settings/Data
    - MRS at 90 and FS at 4 degree in all B-pol
    - Runs 4257, 4260
    
    Progress/positive news
    - T1, T2 seem to work nicely at 4degree 
    - ZDC problems ("negative" pedestals and so on) are mostly
      fixed (DB,FV,AM)
    - No TPC DC HV trip   
    - Update brntuple SummaryPlots (JHL)
      (centrality trigger selections, zdc vs tile)
    
    Problems
    - dead channels in T1: JIJ looking into the problems
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Aug 17 2001 - 16:47:20 EDT